The product of two consecutive even integers is 224. Find the integers
Answer by checkley77(12844)

x(x+2)=224
x^2+2x=224
x^2+2x-224=0
(x+16)(x-14)=0
x+16=0
x=-16 ans.
x-14=0
x=14 ans.